If I have 21 liters of gas held at a pressure of 78 atm and a temperature of 900 K, what will be the volume of the gas if I decrease the pressure to 45 psi and decrease the temperature to 750 K?


1
Expert's answer
2022-03-12T01:59:43-0500

"P1V1T2= P2V2T1"

78× 21×750= 3.06× V2× 900

V2= 1228500/2754

V2= 446L